OpenStudy (anonymous):

write an equation that passes through these points (-2,-2) ,(5,-5) plz help me

OpenStudy (anonymous):

y2-y1/x2-x1 so its -3/7

OpenStudy (anonymous):

(y-y1)/(y2-y1)=(x-x1)/(x2-x1) use this formula

OpenStudy (anonymous):

then u gotta do the other formula which is y-y1=m(x-x1)

OpenStudy (anonymous):

the answer is y= -3/7x-20/7

OpenStudy (anonymous):

did u get it?